Part B Calculate the standard cell potential (E) for the reaction X(s)Y(a)-X (a)+ Y(s) f K-3.70x 104 Express your answer to three significant figures and include the appropriate units View Available Hint(s) Value V SubmitPrev Previous Answers X Incorrect, Try Again; 4 attempts remaining
Answer #1

1.

K = 3.70 * 10-4

E0cell = ( 0.05916 / n ) * LogK

E0cell = ( 0.05916 / 2 ) * Log(3.70*10-4)

E0cell = - 0.102 V

2.

deltaG = deltaH - T deltaS = - 583 - 298.15 ( - 0.257 ) = - 506.4 kJ

3.

E0cell = E0Cu2+/Cu - E0Ni2+/Ni = 0.34 - ( - 0.23 ) = 0.57 V

(0.05916 / n )*logK = E0cell

LogK = 0.57 * 2 / 0.05916

K = 1.86 * 1019

4.

Q = [Co2+]2[Cl2] / [Co3+]2[Cl-]2

Q = ( 0.4142 * 0300 ) / ( 0.2102 * 0.8562 )

Q = 1591.2

Nernst equation,

Ecell = E0cell - ( 0.05916 / n ) * log(Q)

Ecell = 0.483 - ( 0.05916 / 2 ) * log(1591.2)

Ecell = 0.388 V